2. The economy of Newland is in short-run macroeconomic equilibrium. The current real output is $400 billion, and the full employment output is $500 billion. The marginal propensity to consume is 0.8.

(c) Which fiscal policy action, changing government spending or changing taxes, is more effective in closing the output gap? Explain.

(d) Assume instead Newland’s government decides not to take any policy action. Will short-run aggregate supply increase, decrease, or stay the same in the long run? Explain.
1
Expert's answer
2020-03-20T09:48:13-0400

Potential GDP=500, Ep

Actual GDP=400, Ef

Potential GDP>Actual GDP

The economy is experiencing a recession output gap

"Y=(Ep-Ef)\times\Mu=(500-400)\times5=500"

need to increase total output

"MPC+MPS=1"

MPC=0.8, MPS=1-0.8=0.2

"M=\frac{1}{MPS}=\frac{1}{0.2}=5"

"Y=-T\times\frac{MPC}{1-MPC}"

"500=-T\times\frac{0.8}{1-0.8}=-T\times4"

T=-125

need to reduce tax deduction

c)There is not enough demand in the economy, so the following measures are necessary:


  • increase in public procurement
  • reduction in tax rates, the abolition of certain taxes
  • increase in government transfers
  • large-scale construction at the expense of state funds


All these measures should contribute to an increase in aggregate demand.

d) Lack of spending will have a depressing effect on the economy. If you don’t take any action, then deflation will occur. Aggregate demand will slowly fall.